OpenStudy (anonymous):

Which set of numbers represents a Pythagorean triple? A. 6, 9, 12 B. 7, 10, 12 C. 16, 18, 25 D. 27, 36, 45

Directrix (directrix):

Square each number in a triple and then determine if the sum of the squares of two numbers is equal to the other. I'll work out the first option to show what I mean.

Directrix (directrix):

A. 6, 9, 12 6² = 36 9² = 81 12² = 144 Does 36 + 81 = 144 ? Does 117 = 144? No, so option A is not a Pythagorean Triple.

Directrix (directrix):

@shaelyn_linden If you test option B, then I will do option C, okay? B. 7, 10, 12

OpenStudy (anonymous):

@directrix 7^2=49 10^2=100 12^2=144 36+81=117 So B isn't the answer.

Directrix (directrix):

My Turn

Directrix (directrix):

C. 16, 18, 25 16² = 256 18² = 324 25² = 625 Does 625 = 324 + 256? Does 625 = 580 No, so this is not a Pythagorean Triple.

OpenStudy (anonymous):

So it would be D?

Directrix (directrix):

@shaelyn_linden Your Turn for D. 27, 36, 45 I would say that this has to be it but let's check. There may be a typo or trick or something going on.

Directrix (directrix):

Yes, it will be D. I am attaching a list of Pythagorean Triples, well some of them. I'll use D to show you how they can help in your work.

Directrix (directrix):

There are infinitely many Pythagorean Triples. I usually remember (3,4,5); (5,12,13), and (8,15, 17). Here's how they work...

Directrix (directrix):

Look at this: (27, 36, 45) = (9*3, 9*4, 9*5) so (27, 36, 45) is a Pythagorean triple. It is a 9 multiple of the (3,4,5) triple.

Directrix (directrix):

So, if you are doing a problem about the sides of a right triangle, you may be able to save time.